The heights of all female college basketball players produce a distribution that is approximately normal with a mean of 67.6 and a standard deviation of 1.8. The tallest 10% of the players have heights above what value?
Answer #1

Solution : Using Normal Distribution

µ = 67.6
σ = 1.8

for top 10% , z = 1.28

tallest 10% of the players have heights above = zσ + µ = 1.28*1.8+67.6 = 69.904 ~ 69.90

Answer : 69.90